OpenStudy (anonymous):

Find a cubic function, in the form f(x)=ax^3+bx^2+cx+d, that has a local maximum value of 4 at -3 and a local minimum value of 0 at 2. f (x) = ax3 + bx2 + cx + d Math to follow

OpenStudy (anonymous):

\[f \prime(x)=3ax^2+2bx+c\] \[f(-3)=4 => -27a+9b-3c+d\] \[f(2)=0 => a+b+c+d\] \[f \prime(-3)=0 =>27a-6b+c\] \[f \prime(2)=0 => 12a+4b+c\]

OpenStudy (anonymous):

\[\left(\begin{matrix}-27a+9b-3c+d=4 \\ -a+b+c+d=0\end{matrix}\right)=-28a+8b-4c=4\] \[\left(\begin{matrix}27a-6b+c=0 \\-12a+4b+c=0\end{matrix}\right)=12a-10b=0\] \[\left(\begin{matrix}-28a+8b-4c=4 \\ +4(12a+4b+c)=0\end{matrix}\right)=\left(\begin{matrix}-28a+8b-4c=4 \\ +48a+16b+4c=0\end{matrix}\right)\] =20a+24b=4 I need to isolate a and b in order to get d, then c. Help...

OpenStudy (chillout):

Why don't you write a in terms of b?\[=20a+24b=4\ \ \rightarrow 5a +6b=1\rightarrow a= \frac{1-6b}{5}\]

OpenStudy (chillout):

You have a system of 2 equations: 12a−10b=0 and 20a +24b=4

OpenStudy (anonymous):

I'm still unclear as to how to solve this

OpenStudy (anonymous):

Which of the following best describes a line? A. The set of all points in a plane that are equidistant from three points B. The set of all points in a plane at a given distance from a given point C. The set of all points in a plane that are equidistant from two points D. The set of all points in a plane that are equidistant from a given point and a given line

OpenStudy (anonymous):

That is not helpfully to me at all

OpenStudy (anonymous):

I'm attempting to solve it in the manner which I've been shown. I'm not getting this